OpenStudy (anonymous):

Can someone please help me with this!? Write an equation for the line in point/slope form and slope/intercept form that has the given condition. 1a. Slope = � and passes through the origin. 1b. Passing through (-2, 4) and (-1, -1). 1c. x-intercept = -5 and y-intercept = -2. Please and thank you!

OpenStudy (anonymous):

first one is \(y=�x\)

OpenStudy (anonymous):

2nd one Find the Slope First: \[m = \frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

then use oint Slope Form: \[y-y1=m(x-x1)\]

OpenStudy (anonymous):

1c. Use Intercept Form : \[\frac{ x }{ a }+\frac{ y }{ b } = 1\]

OpenStudy (anonymous):

a= x intercept b= y intercept

OpenStudy (anonymous):

So these are 3 seperate equations, right? 1A) What's the slope? -------------------- 1B) Use Point-Slope form. y - y1 = m(x - x1) (-2, 4) and (-1, -1) Find slope: \[m = \frac{ y2 - y1 }{ x2 - x1 } = \frac{ (-1) - 4 }{ (-1) - (-2) } = \frac{ (-1) - 4 }{ -1 + 2 }=\frac{ -5 }{ 1 }= -5\]Slope(m) is -5. So lets use (-2, 4) and the slope (-5) and make an equation in point-slope form. y - y1 = m(x - x1) y - 4 = -5(x -(-2)) y - 4 = -5(x + 2) Get that into Slope-intercept form(y = mx + b), and solve for y. y - 4 = -5(x + 2) Distribute. y - 4 = -5x - 10 Add 4. y = -5x - 6 SO: Point-Slope form: y - 4 = -5(x + 2) Slope-intercept form: y = -5x - 6 ---------------------------------- 1C) x-intercept is -5 y-intercept is -2 SO:  So you have 2 points: (-5, 0) and (0, -2) Find their slope using the same formula as above... \[m = \frac{ y2 - y1 }{ x2 - x1 } = \frac{ (-2) - 0 }{ 0 - (-5) } = \frac{ (-2) - 0 }{ 0 + 5 }=\frac{ -2 }{ 5 }\]Slope(m) is -2/5 Use point (-5,0) and the slope(-2/5) to make the equation. y - 0 = -2/5(x - (-5)) y - 0 = -2/5(x + 5) There is point slope form^ Now slope intercept form. y - 0 = -2/5(x + 5) distribute. y - 0 = -2/5x - 2 add 0. y = -2/5x - 2 SO: Point-Slope form: y - 0 = -2/5(x + 5) Slope-intercept form: y = -2/5x - 2